OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1?

OpenStudy (saifoo.khan):

parallel to that line means same slope. which is -4.

OpenStudy (saifoo.khan):

\[y = -4 x+10\]

OpenStudy (anonymous):

\[(y-(-2))=-4(x-(3))\] \[y+2=-4x+12\] \[y=-4x+10\]

OpenStudy (anonymous):

x + 4y = 10 4x + y = 10 x – 4y = –10 4x – y = –10

OpenStudy (saifoo.khan):

D.